The manager of Moxa is also considering running a full-sized ad in a local newspaper, the Washington Heights News. Running an ad costs $5,000 and 10,000 students read the paper everyday. Of the students who read the paper, 1/2 of them actually notice the ad. Of those who notice the ad, 75% drink coffee. Of those students who notice the advertisement and drink coffee, 20% make the trip to Moxa and buy something. Would handing out fliers be more or less affordable than running an ad in the Washington Heights News? (Not a yes or no answer, explain how you know with numbers)
Answer #1

Total Cost of Running an ad = $5000

Calculating the no of Student who will notice the advertisement and drink coffee & will make trip to Moxa :

No. of Student who read newspaper = 10000 students

No. of students who notice the ad =10000*1/2

= 5000 students

No of Student who will notice the ad drink coffee = 5000*75%

= 3750 students

No of Student who will notice the ad drink coffee & make trip to Moxa = 3750*20%

= 750 students

So, per customer cost of running ad= Total Cost/No. of students who will make trip to Moxa

= 5000/750

= $ 6.6666

So, The cost per Customer for running Washington heights ad is $ 6.666

Since, you have must the handling out fliers part answer, just compare the per customer cost of both.

Note- If you need any clarification regarding this solution, then you can ask in comments

If you like my answer then please Up-vote as it will be motivating.